let f(x)=(2x-pi)^3 + 2x - cosx


then vlue of d\dx(f^-1(x)) at x=pi is?

Grade:

2 Answers

Pramod J AskiitiansExpert-IIT-B
13 years ago

Dear student,

d/dx ( 1/f(x) = -f1(x)/f(x)2

f1(x) = 6*(2x-pi)^2 + 2 +sinx

f1(pi) = 6pi2+2

f(pi) = pi^3 + 2pi + 1

d/dx ( 1/f(x) = -f1(x)/f(x)2 at x= pi => -(6pi2+2)/( pi^3 + 2pi + 1)^2

i hope this explains

ANUVRATHA NARASIMHAN
12 years ago

but here...f^-1 means inverse function of f.........then what would be the answer?